Let f(x) be the recurrence relation defined by fn=fn-12+nfn-2 for n≥2 f0=3 f1=-1 Find f(3)

  1. Let f(x) be the recurrence relation defined by

fn=fn-12+nfn-2 for n≥2

f0=3

f1=-1

Find f(3)
Answer #1

(a) fin) = fort an fra for nya if n=2= fa = f + 2 (2) fo fo= 3; f =- fo- 61) + 4 (3) fe= 11 if n=3 f3 = fat 2(3) f 11+6 (1) =